OpenStudy (anonymous):

Which of the following functions has the largest value when x = 5? p(x) = 2x2 + 4x + 11 h(x) = 4x s(x) = 10x p(x) h(x) s(x) All the functions are equal at x = 5.

OpenStudy (anonymous):

\[p(5) = 2.5^{2} + 4.5 + 11 = 81\]

Mehek (mehek14):

plug 5 for x in each equation and solve \(\bf\color{blue}{p(x) = 2*5^2 + 4*5 + 11}\)

OpenStudy (teddyiswatshecallsme):

p(x) I'm assuming 2x2 is squared which would make p(x) 2*5^2 + 4*5 + 11 2*5 =10^2 = 100 100+ 20 + 11 = 131

OpenStudy (teddyiswatshecallsme):

p(5) = 131

Mehek (mehek14):

\(\bf\color{red}{h(x) = 4*5}\)

OpenStudy (anonymous):

OK thanks you all

Mehek (mehek14):

\(\bf\color{green}{s(x) = 10*5}\)

OpenStudy (anonymous):

P(x) has the largest ending value...... so its A?

Mehek (mehek14):

\(\bf\color{blue}{p(x) = 2*5^2 + 4*5 + 11}\\5^2=25\\2*25=50\\50+4*5+11\\4*5=20\\50+20=70\\70+11=81\)